Calculate dzy dx2 y = -x + et dey dx2 =Calculate dzy dx2 y = 7 x2 day dx2 =The graph of a function is given. Find the approximate coordinates of all points of inflection of the function (if any). (If
Answer #1

a Y = ēn + ax as d lean = a é da 2 1 differentiate wort a dy dn differentiate wohit x day tee da² - е. Y= -1 - 702 x2 . dif